A rich uncle wants to make you a millionaire how much money must he deposit in a trust fund paying 12% compounded quarterly at the time of your birth to yield $1 million when you retire at age 58 round your answer to the nearest cent
1
Expert's answer
2020-11-02T18:54:07-0500

If one fixes the initial balance "P_0" , the nominal interest rate "r" and the duration of the deposit "T" (in years), you earn more interest with more compounding periods per year "K". The number of compounding periods that make up "T" will be "KT". Then we use the formula

"P_{KT}=P_0(1+\frac{r}{K})^{KT}"

In our case, "r=0.12, K=4, KT=58\cdot 4=232, P_{KT}=1,000,000".


Therefore, we have that "1,000,000=P_0(1+\frac{0.12}{4})^{232}=P_0(1.03)^{232}=951.12179\cdot P_0", and consequently, "P_0=\frac{1,000,000}{951.12179}=1,051.39".


Answer: a rich uncle must deposit in a trust fund $1,015.39 to make you a millionaire when you retire at age 58.